OpenStudy (anonymous):

PLZ I has to finish this up now Which of the following expressions are equivalent to −3.5(2 − 3n) − 2.5n? Select all that apply. A. −7 − 8n B. −7 + 8n C. −7 − 13n D. −7 − n(10.5 − 2.5) E. −7 + n(10.5 − 2.5)

OpenStudy (anonymous):

Pwease halp

OpenStudy (anonymous):

−3.5(2 − 3n) − 2.5n = -7 +10.5n - 2.5n

OpenStudy (anonymous):

-3.5(2-3n) -2.5n= -7 +10.5n - 2.5n

OpenStudy (anonymous):

B. and E.

OpenStudy (anonymous):

Hmmm intresting

OpenStudy (anonymous):

Can I ask another question?

OpenStudy (anonymous):

NVM